OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it diverges, find the limit: an=2^(-n)cosnpi

OpenStudy (anonymous):

\(\large a_{n}=2^{-n} \cos n \pi = \frac{(-1)^{n}}{2^n} \)

OpenStudy (anonymous):

'math processing error' thats not what i wrote ! lol

OpenStudy (anonymous):

hahaha, no, i got what you typed. how come you ignored the cosnπ?

OpenStudy (anonymous):

@nuk64 , we cant find a limit if the series diverges. limit can only be found for converging sequences

OpenStudy (anonymous):

@vamgadu : How do we know this one's diverging?

OpenStudy (anonymous):

see